# The Quantum Esscher Transform

### v1

@@ 4.3 Further discussion
If the desired output model is a normalized state, one can apply similar techniques for Gibbs sampling to extract the normalized Esscher-transformed state from the output of Algorithm 1. We briefly describe this procedure and the overhead cost it incurs. More details can be found in Chapter 3 of [Gil19]. Let $\varepsilon>0$ denote the desired precision in trace distance between our approximate output and the ideal state. First, we prepare a maximally entangled state on two registers. Use Algorithm 1 to construct a $1$-block-encoding $U$ of $e^{\frac{\sum_{i}\theta_{i}H_{i}+\log\rho}{2}}/\sqrt{\mathcal{N}}$ where $\mathcal{N}=e^{\|\theta\|_{1}+2(1+\log 2\kappa)}$, with block-encoding error $0<\varepsilon_{1}<\varepsilon/N^{2}$. Then apply $U$ to the second register to obtain a state $|\psi\rangle$, so that tracing out the first register yields an approximate subnormalized state with trace distance error of $\mathcal{O}\left(\varepsilon/N\right)$. That is,
@@
With $\mathcal{Z}:=\operatorname{Tr}\left(e^{\sum_{i}\theta_{i}H_{i}+\log\rho}\right)$, this state, when postselected after $\mathcal{O}\left(\sqrt{\frac{N\mathcal{N}}{\mathcal{Z}}}\log\frac{1}{\varepsilon}\right)$ steps of fixed-point amplitude amplification (refer to Theorem 27 in [GSLW19]), results in a density operator $\varepsilon$-close to the normalized Esscher-transformed state
@@
$\widetilde{\mathcal{O}}\left(\kappa\log^{2}\left(\frac{N^{2}}{\varepsilon}\right)\right)\cdot\mathcal{O}\left(\sqrt{\frac{N\mathcal{N}}{\mathcal{Z}}}\log\frac{1}{\varepsilon}\right)\subseteq\widetilde{\mathcal{O}}\left(\kappa\sqrt{\frac{N\mathcal{N}}{\mathcal{Z}}}\log^{3}\left(\frac{1}{\varepsilon}\right)\right).$
5 Conclusion
@@
  • Is there a quantum algorithmic framework that can fully solve the minimum relative entropy problem? Our current approach only presents the formal solution for the optimal parameter $\lambda^{*}$. Approaches such as Newton’s algorithm with backtracking was suggested in [ZTF13], the quantized version of which could be studied. Additionally, [AAKS20] demonstrated that $\lambda^{*}$ can, in principle, be found with a convex optimization program. Can we design a quantum algorithm to effectively address this problem?
  • One could explore strategies for alternative input models. Our current work exclusively considered the purified access model, wherein the preparation of the purification of the input state was assumed. In contrast, the sampling access model, which assumes multiple independent copies of the input state, is another commonly used model. Gilyén et al. [GP22] has proposed an approach to implement approximate block-encodings of $\rho$, starting with sample access. This approach is based on a combination of density matrix exponentiation [LMR14, ${}^{+}$17] and QSVT, and allows us to implement the quantum Esscher transform in the sampling access model. We leave the total cost of this procedure for further analysis.
  • In Section 2.2.3, we noted potential connections between the quantum Esscher transform and imaginary-time evolution. To give these substance, further investigation is required.
@@ Acknowledgments
The authors would like to thank Po-Wei Huang, Xiufan Li, Zhan Yu, Serge Massar and Roberto Rubboli for helpful discussions. This work is supported by the National Research Foundation, Singapore, and A*STAR under its CQT Bridging Grant and its Quantum Engineering Programme under grant NRF2021-QEP2-02-P05. KK acknowledges support from Leong Chuan Kwek, under project grant R-710-000-007-135.
Appendix A Proof of Theorem 2.2
@@
Lemma A.1.
(Proposition 3.24 – [FS11]) Let $P$ be a probability measure on $(\Omega,\Sigma)$ and $X$ be a random variable on $\Omega$. Fix $m\in\mathbbm{R}^{d}$. Then for any probability measure $Q$ on $(\Omega,\Sigma)$ satisfying $\mathbb{E}_{Q}[X]=m$, we have
@@
[…]
@@
Since this holds for all $\lambda\in\mathbb{R}^{d}$, we conclude that $D(Q\|P)\geq\sup_{\lambda\in\mathbb{R}^{d}}\left[\lambda\cdot m-\log\mathbb{E}_{P}[e^{\lambda\cdot X}]\right]$. Furthermore, if $\lambda^{\prime}\in\mathbb{R}^{d}$ is such that $Q_{\lambda^{\prime}}\in\Lambda\cap M$, then letting $Q=Q_{\lambda^{\prime}}$ and rerunning the same argument sequence above gives
@@
The last equivalence holds because $\log f(x)$ and $f(x)$ share the same minimum/maximum points, provided $f(x)>0$ at those points. It remains to show $Q^{\star}$ indeed minimizes $D(Q\|P)$, subject to the constraints $E_{Q}[X]=m$. But this follows easily from Lemma A.1. Furthermore, since $x\mapsto x\log x$ is a strictly convex function, $D(Q\|P)$ is a strictly convex functional of $Q$ and so it can have at most one minimizer in the convex set $M$, thereby showing the uniqueness of $Q^{\star}$. Finally, again using Lemma A.1 we have $\lambda^{\star}=\operatorname*{argmax}_{\lambda\in\mathbb{R}^{d}}\left[\lambda\cdot m-\log\mathbb{E}_{P}[e^{\lambda\cdot X}]\right]=\operatorname*{argmin}_{\lambda\in\mathbb{R}^{d}}\left[\log\mathbb{E}_{P}[e^{\lambda\cdot(X-m)}]\right]=\operatorname*{argmin}_{\lambda\in\mathbb{R}^{d}}\mathbb{E}_{P}[e^{\lambda\cdot(X-m)}]$. ∎
@@ Appendix B Wirtinger Calculus
The ‘Wirtinger Calculus’ provides a methodology for optimization problems involving complex matrices. It enables ‘differentiation as usual’ with respect to complex matrices. In this appendix, we state only the main definitions and results needed to solve Problem 2.4. For a more thorough exposition of this framework, we direct the reader to [KQKR23, Hjø11, KD09].
@@
The following three propositions are all we need in this paper. We omit their proofs, which can all be found in [KQKR23].
Proposition B.1.
@@
Proposition B.3 (Wirtinger derivatives with respect to Hermitian matrices).

### v2

@@ III.3 Further discussion
If the desired output model is a normalized state, one can apply similar techniques for Gibbs sampling to extract the normalized Esscher-Transformed state from the output of Algorithm 1. We briefly describe this procedure and the overhead cost it incurs. More details can be found in Chapter 3 of [10]. Let $\varepsilon>0$ denote the desired precision in trace distance between our approximate output and the ideal state. First, we prepare a maximally entangled state on two registers. Use Algorithm 1 to construct a $1$-block-encoding $U$ of $e^{\frac{\sum_{i}\theta_{i}H_{i}+\log\rho}{2}}/\sqrt{\mathcal{N}}$ where $\mathcal{N}=e^{\|\theta\|_{1}+2(1+\log 2\kappa)}$, with block-encoding error $0<\varepsilon_{1}<\varepsilon/N^{2}$. Then apply $U$ to the second register to obtain a state $|\psi\rangle$, so that tracing out the first register yields an approximate subnormalized state with trace distance error of $\mathcal{O}\left(\varepsilon/N\right)$. That is,
@@
With $\mathcal{Z}:=\operatorname{Tr}\left(e^{\sum_{i}\theta_{i}H_{i}+\log\rho}\right)$, this state, when postselected after $\mathcal{O}\left(\sqrt{\frac{N\mathcal{N}}{\mathcal{Z}}}\log\frac{1}{\varepsilon}\right)$ steps of fixed-point amplitude amplification (refer to Theorem 27 in [15]), results in a density operator $\varepsilon$-close to the normalized Esscher-Transformed state
@@
$\widetilde{\mathcal{O}}\left(\kappa\log^{2}\left(\frac{N^{2}}{\varepsilon}\right)\right)\cdot\mathcal{O}\left(\sqrt{\frac{N\mathcal{N}}{\mathcal{Z}}}\log\frac{1}{\varepsilon}\right)$
$\subseteq\widetilde{\mathcal{O}}\left(\kappa\sqrt{\frac{N\mathcal{N}}{\mathcal{Z}}}\log^{3}\left(\frac{1}{\varepsilon}\right)\right).$
IV Conclusion
@@
  • Is there a quantum algorithmic framework that can fully solve the minimum relative entropy problem? Our current approach only presents the formal solution for the optimal parameter $\lambda^{*}$. Approaches such as Newton’s algorithm with backtracking was suggested in [38], the quantized version of which could be studied. Additionally, [1] demonstrated that $\lambda^{*}$ can, in principle, be found with a convex optimization program. Can we design a quantum algorithm to effectively address this problem?
  • One could explore strategies for alternative input models. Our current work exclusively considered the purified access model, wherein the preparation of the purification of the input state was assumed. In contrast, the sampling access model, which assumes multiple independent copies of the input state, is another commonly used model. Gilyén et al. [13] has proposed an approach to implement approximate block-encodings of $\rho$, starting with sample access. This approach is based on a combination of density matrix exponentiation [24, 20] and QSVT, and allows us to implement the quantum Esscher Transform in the sampling access model. We leave the total cost of this procedure for further analysis.
  • In Section II.2, we noted potential connections between the quantum Esscher Transform and imaginary-time evolution. To give these substance, further investigation is required.
@@ Acknowledgments
The authors would like to thank Po-Wei Huang, Xiufan Li, Zhan Yu, Roberto Rubboli and Serge Massar for helpful discussions. This work is supported by the National Research Foundation, Singapore, and A*STAR under its CQT Bridging Grant and its Quantum Engineering Programme under grant NRF2021-QEP2-02-P05. KK acknowledges support from Leong Chuan Kwek, under project grant R-710-000-007-135.
Appendix A Proof of Theorem 1
@@
Lemma 6.
(Proposition 3.24 – [9]) Let $P$ be a probability measure on $(\Omega,\Sigma)$ and $X$ be a random variable on $\Omega$. Fix $m\in\mathbbm{R}^{d}$. Then for any probability measure $Q$ on $(\Omega,\Sigma)$ satisfying $\mathbb{E}_{Q}[X]=m$, we have
@@
$\geq$ $\sum_{\omega\in\Omega}Q(\omega)\log\frac{Q_{\lambda}(\omega)}{P(\omega)}$
@@
The third inequality is due to Jensen’s inequality $D(Q\|P)\geq 0$. Since this holds for all $\lambda\in\mathbb{R}^{d}$, we conclude that $D(Q\|P)\geq\sup_{\lambda\in\mathbb{R}^{d}}\left[\lambda\cdot m-\log\mathbb{E}_{P}[e^{\lambda\cdot X}]\right]$. Furthermore, if $\lambda^{\prime}\in\mathbb{R}^{d}$ is such that $Q_{\lambda^{\prime}}\in\Lambda\cap M$, then letting $Q=Q_{\lambda^{\prime}}$ and rerunning the same argument sequence above gives
@@
The last equivalence holds because $\log f(x)$ and $f(x)$ share the same minimum/maximum points, provided $f(x)>0$ at those points. It remains to show $Q^{\star}$ indeed minimizes $D(Q\|P)$, subject to the constraints $E_{Q}[X]=m$. But this follows easily from Lemma 6. Furthermore, since $x\mapsto x\log x$ is a strictly convex function, $D(Q\|P)$ is a strictly convex functional of $Q$ and so it can have at most one minimizer in the convex set $M$, thereby showing the uniqueness of $Q^{\star}$. Finally, again using Lemma 6 we have $\lambda^{\star}=\operatorname*{argmax}_{\lambda\in\mathbb{R}^{d}}\left[\lambda\cdot m-\log\mathbb{E}_{P}[e^{\lambda\cdot X}]\right]=\operatorname*{argmin}_{\lambda\in\mathbb{R}^{d}}\left[\log\mathbb{E}_{P}[e^{\lambda\cdot(X-m)}]\right]=\operatorname*{argmin}_{\lambda\in\mathbb{R}^{d}}\mathbb{E}_{P}[e^{\lambda\cdot(X-m)}]$. ∎
Remark 2.
The random variable $X$ induces from the probability measure $P$ the probability mass function $P_{X}(x):=P(X^{-1}(x))$ on $E:=X(\Omega)$. Assume we have, for probability measures $Q,P$ and random variable $X$, that
Then for the probability mass functions $Q_{X}$ and $P_{X}$ we have
$Q_{X}(x)=Q(X^{-1}(x))$ $=\sum_{\omega:X(\omega)=x}Q(\omega)$
$=\frac{e^{\theta\cdot x}P_{X}(x)}{\sum_{x\in E}e^{\theta\cdot x}P_{X}(x)},$
i.e., $Q_{X}$ is the Esscher Transform of $P_{X}$ as defined above.
@@ Appendix B Wirtinger Calculus
The ‘Wirtinger Calculus’ provides a methodology for optimization problems involving complex matrices. It enables ‘differentiation as usual’ with respect to complex matrices. In this appendix, we state only the main definitions and results needed to solve Problem 1. For a more thorough exposition of this framework, we direct the reader to [21, 16, 19].
@@
The following three propositions are all we need in this paper. We omit their proofs, which can all be found in [21].
Proposition 7.
@@
Proposition 9 (Wirtinger derivatives with respect to Hermitian matrices).
@@
Appendix C Proof of Theorem 2
Proof.
Step 1. First, for any candidate solution $\sigma$ we enforce $\ker\rho\subseteq\ker\sigma$. By Lemma 10, this implies $\sigma_{\ker\rho}=\mathbf{0}$ and furthermore enables the decomposition of $\sigma$ into a direct sum: $\sigma=\sigma_{\operatorname{supp}\rho}\oplus\sigma_{\ker\rho}$. With this decomposition, we can consider the trace of the operators over just the subspace $\operatorname{supp}\rho$. More specifically, $\operatorname{Tr}(\sigma H_{i})=\operatorname{Tr}(\sigma(\Pi_{\operatorname{supp}\rho}+\Pi_{\ker\rho})H_{i}(\Pi_{\operatorname{supp}\rho}+\Pi_{\ker\rho}))=\operatorname{Tr}(\sigma_{\operatorname{supp}\rho}H_{i,\operatorname{supp}\rho})$ [^2] Recall that for any $A\in\mathcal{L}(\mathcal{H})$, $\ker A\oplus\operatorname{supp}A=\mathcal{H}$, so $\Pi_{\ker A}+\Pi_{\operatorname{supp}A}=I$. and
$S(\sigma\|\rho)$ $=$ $\operatorname{Tr}\{\sigma_{\operatorname{supp}\rho}\oplus\sigma_{\ker\rho}\left(\log(\sigma_{\operatorname{supp}\rho}\oplus\sigma_{\ker\rho})-\log(\rho_{\operatorname{supp}\rho}\oplus\rho_{\ker\rho})\right)\}$ (C.1)
$=$ $\operatorname{Tr}\{\sigma_{\operatorname{supp}\rho}(\log\sigma_{\operatorname{supp}\rho}-\log\rho_{\operatorname{supp}\rho})\}+\underbrace{\operatorname{Tr}\{\sigma_{\ker\rho}(\log\sigma_{\ker\rho}-\log\rho_{\ker\rho})\}}_{=0}$ (C.2)
$=$ $S(\sigma_{\operatorname{supp}\rho}\|\rho_{\operatorname{supp}\rho}).$ (C.3)
Thus, we can replace $\mathcal{H}$ in Problem 1 by $\operatorname{supp}\rho$, and the operators by their restrictions to $\operatorname{supp}\rho$. Note that $\rho_{\operatorname{supp}\rho}$ is positive definite.
Step 2. Next we obtain the form of $\sigma_{\operatorname{supp}\rho}$. For ease of presentation let us simply denote $(\sigma/\rho/H_{i})_{\operatorname{supp}\rho}$ by $(\sigma/\rho/H_{i})$. With $\rho$ now positive definite, $\log\rho$ is well-defined. Now we invoke Proposition 7 to extract the optimal $\sigma$ by setting $\frac{\partial\mathcal{L}}{\partial\sigma}=\mathbf{0}$. Set up the Lagrangian
$\mathcal{L}=\operatorname{Tr}\{\sigma(\log\sigma-\log\rho)\}-\sum_{i}\lambda_{i}(\operatorname{Tr}(\sigma H_{i})-m_{i})-\eta(\operatorname{Tr}\sigma-1)$ (C.4)
where $\lambda_{i}$ and $\eta$ are the Lagrange multipliers. Making use of Propositions 8 and 9, setting $\frac{\partial\mathcal{L}}{\partial\sigma}$ to zero gives
$\implies\sigma=e^{\eta-1}e^{\lambda\cdot H+\log\rho}$
Step 3. Now we show that $\sigma^{\star}$ as given in Eq. II.4 indeed minimizes $S(\sigma\|\rho)$. But this follows easily from Lemma 11. Furthermore, since $S(\sigma\|\rho)$ is a strictly convex functional of $\sigma$, it can have at most one minimizer in the convex set $M$, thereby showing the uniqueness of $\sigma^{\star}$. Finally, again by Lemma 11 we note that $\lambda^{\star}$ satisfies $\lambda^{\star}=\operatorname*{argmax}_{\lambda\in\mathbb{R}^{d}}\left[\lambda\cdot m-\log\operatorname{Tr}(e^{\lambda\cdot H+\log\rho})\right]=\operatorname*{argmin}_{\lambda\in\mathbb{R}^{d}}\log\operatorname{Tr}(e^{\lambda\cdot(H-m)+\log\rho})=\operatorname*{argmin}_{\lambda\in\mathbb{R}^{d}}\operatorname{Tr}(e^{\lambda\cdot(H-m)+\log\rho})$, where the last equality holds because $\log f(x)$ and $f(x)$ share the same minimum/maximum points, provided $f(x)>0$ at those points. ∎
Lemma 10.
$\sigma=\sigma_{\operatorname{supp}\rho}\oplus\sigma_{\ker\rho}.$
Proof.
$\sigma=$ $\sum_{i,j=0}^{N-1}\braket{i}{\sigma}{j}\ket{i}\bra{j}$ |
[…]
The following lemma is the quantized version of Lemma 6. We employ analogous arguments and notation, starting with
Lemma 11.
$S(\sigma\|\rho)\geq\sup_{\lambda\in\mathbb{R}^{d}}\left[\lambda\cdot m-\log\operatorname{Tr}(e^{\lambda\cdot H+\log\rho})\right].$ (C.5)
$S(\sigma_{\lambda^{\prime}}\|\rho)=\lambda^{\prime}\cdot m-\log\operatorname{Tr}(e^{\lambda^{\prime}\cdot H+\log\rho})=\sup_{\lambda\in\mathbb{R}^{d}}\left[\lambda\cdot m-\log\operatorname{Tr}(e^{\lambda\cdot H+\log\rho})\right].$ (C.6)
Proof.
$S(\sigma\|\rho)$ $=$ $S(\sigma\|\sigma_{\lambda})+\operatorname{Tr}\{\sigma(\log\sigma_{\lambda}-\log\rho)\}$ (C.7)
Appendix D Block-encodings and Quantum Singular Value Transformation
The technique of quantum signal processing [25] and its lifting via ‘qubitization’ to quantum singular value transformation (QSVT) [23, 15] provide a concise way to formulate quantum algorithms, particularly for linear algebraic tasks. This framework has provided more efficient implementations of several existing quantum algorithms, such as Hamiltonian simulation [22, 23], amplitude amplification and estimation [15, 33] and quantum linear systems solving [15], and even led to the discovery of new algorithms. For our purposes, we do not actually need the full generality of QSVT. As our matrices of interest are Hermitian and thus admit spectral decompositions, a relaxed version of QSVT—quantum eigenvalue transformation (QET)—suffices. We direct readers interested in learning more about QSVT to [15, 27, 7].
Definition 3 (Block-Encoding).
Remark 3.
[15] provides a construction of exact block-encodings for density operators, assuming access to oracles which prepare the purifications of the density operators:
Definition 4 (Purified quantum query-access).
[…]
Proposition 12 (Block-encoding of density operators – Lemma 45, [15]).
For general matrices which need not be density operators, [6, 15] also showed how to implement their block-encodings efficiently, assuming the existence of quantum random access memory (QRAM) [12]. Given block-encodings of operators $A_{i}$, we can construct block-encodings of their linear combinations and products. For linear combinations, we make use of an auxiliary tool known as a ‘state preparation pair’. Recall that $\|\cdot\|_{1}$ is the $l_{1}$/Manhattan norm.
Definition 5 (State Preparation Pair).
[…]
Proposition 13 (Linear combination of block-encoded matrices – Lemma 52, [15]).
Let
  1. i.
  2. ii.
    $A=\sum_{j=0}^{m-1}y_{j}A_{j}$ for $y:=(y_{0},\dots,y_{m-1})\in\mathbb{C}^{m}$,
  3. iii.
where
[…]
is a $(n+a+b)$-qubit unitary.
In Proposition 13, the subnormalization factors of the $A_{j}$’s are restricted to be the same. Later on, we will need a slight generalization of the above result whereby this requirement is dropped.
Proposition 14 (Generalized linear combination of block-encoded matrices).
Let
  1. i.
  2. ii.
    $A=\sum_{j=0}^{m-1}y_{j}A_{j}$ for $y:=(y_{0},\dots,y_{m-1})\in\mathbb{C}^{m}$,
  3. iii.
where
[…]
is a $(n+a+b)$-qubit unitary.
Proof.
The following is adapted from the proof of Lemma 52, [15]. By definition of state-preparation pairs (see Definition 5), $P_{L}\ket{0^{b}}=\sum_{j=0}^{2^{b}-1}c_{j}\ket{j}$ and $P_{R}\ket{0^{b}}=\sum_{j=0}^{2^{b}-1}d_{j}\ket{j}$ such that $\sum_{j=0}^{m-1}|\alpha_{j}y_{j}-\beta c_{j}^{*}d_{j}|\leq\varepsilon_{\text{SP}}$. First we evaluate the block extraction of $\widetilde{W}$. We have
$\quad\;(\bra{0^{b+a}}\otimes I_{n})\widetilde{W}(\ket{0^{b+a}}\otimes I_{n})$
[…]
$\leq$ $\varepsilon_{\text{SP}}+\sum_{j=0}^{m-1}|y_{j}|\varepsilon_{\text{BE}}\leq\varepsilon_{\text{SP}}+\frac{\beta}{\inf_{j}\alpha_{j}}\varepsilon_{\text{BE}}.$
Remark 4.
In the special case where the block-encodings of the $A_{j}$’s have the same subnormalization factors, i.e., $\alpha_{j}=\alpha$ for all $j$, we recover Proposition 13 from Proposition 14 . To see this, observe that if $(P_{L},P_{R})$ is a $(\beta,b,\varepsilon_{\text{SP}})$-state-preparation-pair for $\alpha\odot y$, then $\sum_{j}|\alpha_{j}y_{j}-\beta c_{j}^{*}d_{j}|\leq\varepsilon_{\text{SP}}\implies\sum_{j}|\alpha y_{j}-\beta c_{j}^{*}d_{j}|\leq\varepsilon_{\text{SP}}\implies\sum_{j}|y_{j}-\frac{\beta}{\alpha}c_{j}^{*}d_{j}|\leq\frac{\varepsilon_{\text{SP}}}{\alpha}$, thus implying $(P_{L},P_{R})$ is a $(\frac{\beta}{\alpha},b,\frac{\varepsilon_{\text{SP}}}{\alpha})$-state-preparation-pair for $y$. According to Proposition 13, $\widetilde{W}$ is then a $(\alpha\cdot\frac{\beta}{\alpha},\;a+b,\;\alpha\cdot\frac{\varepsilon_{\text{SP}}}{\alpha}+\frac{\beta}{\alpha}\varepsilon_{\text{BE}})$-BE of $A$. This is in agreement with Proposition 14.
Theorem 15 (Polynomial Eigenvalue Transformation – Theorem 56, [15]).
Proposition 16 (Bounded Polynomial Approximation – Corollary 66, [15]).
$\|f(x)-P(x)\|_{[x_{0}-r,x_{0}+r]}\leq\varepsilon$ (D.2)
$\|P(x)\|_{[-1,1]}\leq\varepsilon+\|f(x)\|_{[x_{0}-r-\delta/2,x_{0}+r+\delta/2]}\leq\varepsilon+B$ (D.3)
$\|P(x)\|_{[-1,1]\setminus[x_{0}-r-\delta/2,x_{0}+r+\delta/2]}\leq\varepsilon.$ (D.4)
Theorem 15 and Proposition 16 are to be used in conjunction to produce block-encodings of general functions of Hermitian matrices. In doing so, we first note that Theorem 15 produces an encoding of $P(A/\alpha)$, not $P(A)$. Thus, with a polynomial approximation of $f$, say $P(x)\approx f(x)$, it is generally not true that $P(A/\alpha)\approx f(A)$. What we need is a polynomial approximation not of $f$, but of a (horizontally) scaled version of $f$, $f^{\prime}(x):=f(\alpha x)$, so that $P(x)\approx f^{\prime}(x)\implies P(A/\alpha)\approx f^{\prime}(A/\alpha)=f(A)$. Second, we also have to take into account the polynomial approximation error incurred in producing the final desired block encoding $f(A)$. We take care of these matters in Corollary 17, which, given the block-encoding of an arbitrary Hermitian matrix $A$, produces a block-encoding of $f(A)$, where $f$ is a generic real-valued function.
Corollary 17 (Block-encoding functions of general Hermitian matrices).
Given
  1. i.
  2. ii.
    $f:I\longrightarrow\mathbb{R}$, a smooth function on an open interval $I$ containing $[\lambda_{\min},\lambda_{\max}]$. Assume the function $x\mapsto f(\alpha x)$ satisfies the conditions in Proposition 16 with $[\lambda_{\min}/\alpha,\lambda_{\max}/\alpha]\subseteq[x_{0}-r,x_{0}+r]$ and series-of-coefficients bound $B$.
  3. iii.
Proof.
By Proposition 16, given polynomial approximation error tolerance $\varepsilon_{\text{poly}}$ there exists a polynomial $Q\in\mathbb{C}[x]$ of degree $\mathcal{O}\left(\frac{1}{\delta}\log\left(\frac{B}{\varepsilon_{\text{poly}}}\right)\right)$ which $\varepsilon_{\text{poly}}$-approximates $f\circ t_{\alpha}^{-1}$ on $[x_{0}-r,x_{0}+r]$ and is bounded above by $1+B$ on $[-1,1]$. Since $\|A/\alpha\|\in[\lambda_{\min}/\alpha,\lambda_{\max}/\alpha]\subseteq[x_{0}-r,x_{0}+r]$, we have
In order to apply Theorem 15, our polynomial has to be real and upper-bounded by $1/2$ on $[-1,1]$. Observe that for any complex-valued function $F$ and domain $S$,
Since $f$ itself is real-valued, $\operatorname{Re}Q\in\mathbb{R}[x]$ is qualified to assume the role of $P$ in Proposition 16. That is, the real polynomial $\operatorname{Re}Q$ also $\varepsilon_{\text{poly}}$-approximates $f\circ t_{\alpha}^{-1}$ on $[x_{0}-r,x_{0}+r]$ and is bounded above by $1+B$ on $[-1,1]$. Thus, letting $P\leftarrow\frac{\operatorname{Re}Q}{2(1+B)}$ in Theorem 15 we obtain $\tilde{U}$, a $(1,a+2,4d\sqrt{\varepsilon/\alpha})$-encoding of $\frac{\operatorname{Re}Q}{2(1+B)}(A/\alpha)$, where $d=\mathcal{O}\left(\frac{1}{\delta}\log\left(\frac{B}{\varepsilon_{\text{poly}}}\right)\right)$. Putting these together and noting that $f\circ t_{\alpha}^{-1}(\frac{A}{\alpha})=f(A)$, we have
$\left\|\frac{f(A)}{2(1+B)}-(\bra{0^{a+2}}\otimes I)\tilde{U}(\ket{0^{a+2}}\otimes I)\right\|$ $\leq\left\|\frac{f\circ t_{\alpha}^{-1}(\frac{A}{\alpha})}{2(1+B)}-\frac{\operatorname{Re}Q(\frac{A}{\alpha})}{2(1+B)}\right\|+\left\|\frac{\operatorname{Re}Q(\frac{A}{\alpha})}{2(1+B)}-(\bra{0^{a+2}}\otimes I)\tilde{U}(\ket{0^{a+2}}\otimes I)\right\|$
$\leq\frac{\varepsilon_{\text{poly}}}{2(1+B)}+4d\sqrt{\varepsilon/\alpha}.$